# Private Benchmarking to Prevent Contamination and Improve Comparative Evaluation of LLMs

### v1

@@ Abstract
Benchmarking is the de-facto standard for evaluating LLMs, due to its speed, replicability and low cost. However, recent work has pointed out that the majority of the open source benchmarks available today have been contaminated or leaked into LLMs, meaning that LLMs have access to test data during pretraining and/or fine-tuning. This raises serious concerns about the validity of benchmarking studies conducted so far and the future of evaluation using benchmarks. To solve this problem, we propose Private Benchmarking, a solution where test datasets are kept private and models are evaluated without revealing the test data to the model. We describe various scenarios (depending on the trust placed on model owners or dataset owners), and present solutions to avoid data contamination using private benchmarking. For scenarios where the model weights need to be kept private, we describe solutions from confidential computing and cryptography that can aid in private benchmarking. Finally, we present solutions the problem of benchmark dataset auditing, to ensure that private benchmarks are of sufficiently high quality.
Private Benchmarking to Prevent Contamination and Improve Comparative Evaluation of LLMs
Nishanth Chandran Sunayana Sitaram Divya Gupta
Rahul Sharma Kashish Mittal Manohar Swaminathan
Microsoft Research India nichandr, sunayana.sitaram@microsoft.com
@@ 1 Introduction
Large Language Models (LLMs) have become increasingly popular due to their impressive performance in various tasks and domains, often surpassing human capabilities. However, it is crucial for LLMs to not only excel in their tasks but also be grounded in factual information and avoid generating harmful content. Therefore, evaluating and understanding the limitations and capabilities of LLMs has become an important topic of research. Ensuring that LLMs possess these essential qualities is crucial for their widespread adoption in various applications.
LLMs are commonly evaluated through automated techniques like benchmarking, or by human interaction and feedback. Holistic evaluation of LLMs Liang et al. (2022) evaluates LLMs on several dimensions such as accuracy, calibration, bias and toxicity. Several popular benchmarks have been proposed such as BigBench Srivastava et al. (2023) to increase the task and language coverage of benchmarking. Recently, LLM-based evaluation, in which an LLM is used to evaluate the output of another LLM has also become popular. While human evaluation is often considered the gold standard, it can be slow, expensive and difficult to scale. LLM-based evaluation is much faster and may be less expensive than human evaluation, but is known to have biases, inconsitencies and poor agreement with human judgments in some cases Clark et al. (2021); Chen et al. (2023); Wang et al. (2023); Chiang and Lee (2023); Golchin and Surdeanu (2023a).
@@
Benchmarking too has a number of limitations. For instance, the tasks and datasets provided in standard open source benchmarks may not always be applicable in real-world scenarios. Additionally, certain benchmarks may be of low quality due to inadequate or biased instructions Parmar et al. (2023) or limited training of annotators Geva et al. (2019). Moreover, some benchmarks may contain unintended artifacts introduced by annotators’ subconscious biases Poliak et al. (2018). Nevertheless, in spite of these constraints, benchmarking remains a fast, cost-effective, and replicable means of comparing multiple models or versions of a single model. As such, benchmarking has become the de-facto standard for evaluating the capabilities of LLMs, showing superiority of one LLM over another and deciding which version of a model to ship in production.
@@
However, the issue of test dataset contamination or leakage has raised significant concerns about the reliability of benchmarking and the validity of its outcomes. Benchmark datasets are commonly found on the internet, allowing large web crawlers that collect LLM pretraining data to potentially incorporate them into pretraining data. Another concerning issue, as pointed out by Balloccu et al. (2024), is that LLM researchers could have inadvertently leaked benchmark datasets to closed source models during LLM evaluation, and this data can be used for fine-tuning or prompt optimization. According to estimates based on their survey, over 4 million data points have been leaked to closed models such as ChatGPT. Creating benchmarks can be a costly process due to the human effort and expertise required for data creation and annotation. Therefore, it is imperative to prevent contamination of benchmarks.
@@
In this paper, we describe Private Benchmarking - a solution that keeps the benchmark dataset private from the model, while revealing only the results of the evaluation. This ensures that the model never sees the test dataset. There are several possible scenarios - the model may be open source, in which case it can be hosted locally and a private benchmark can be run on it. In case of closed models, we may trust the model owner not to use any benchmark dataset for fine-tuning or model improvements. However, as pointed out by Balloccu et al. (2024), it can be challenging for model owners to identify benchmark data and filter it out to avoid training on it. In case of closed models, the model owner may want to keep models and their weights private. Further, if the benchmark datasets are kept private, how does one guarantee that it has test data points that are of good quality? Ideally, we want solutions where an auditor is able to check a private benchmark for quality. In this work, we present unique solutions to these problems by introducing advances in confidential computing and cryptographic protocols to the NLP benchmarking community. As far as we are aware, this is the first work that has approached the problem of benchmark data contamination with this perspective.
@@ 2 Challenges with LLM benchmarking today
As mentioned earlier, several challenges exist with benchmarking practices today. Benchmarks are expensive to create due to the human effort typically involved in data cleaning, preparation and annotation. The pace at which benchmarks are becoming saturated is increasing Potts (2021), making them less useful much more quickly than before.
Benchmarks, especially those created in the pre-LLM era are created on tasks that do not have real-world equivalents, making it difficult to interpret the value of improvements on the benchmark. Examples of such benchmarks include those created for canonical NLP tasks such as Part of Speech tagging and Sentiment Analysis Hu et al. (2020). Major developers of LLMs also continuously create benchmarks to internally evaluate and iteratively improve their models. Creating such benchmarks requires considerable expertise and time. Such internal benchmarks also hold a competitive advantage to the model owner, as their secret sauce that results in their models outperforming others. Hence model builders in the industry do not always share their benchmarks. Several commercial organizations building applications on top of LLMs also utilize their own internal benchmarks based on user data to evaluate their own models before shipping to production - however, such datasets are not typically made available for research use due to concerns about privacy and sensitivity of the data. If there are mechanisms for competing models to be evaluated on competing benchmarks with neither the model details nor the benchmarks data set being revealed to either party, it will greatly enhance credible benchmarking and faster improvements of models themselves. Access to more benchmarks created by industry will also boost LLM and evaluation research, particularly in the academic community.
@@ 2.1 Data contamination in LLMs
In the context of Large Language Models (LLMs), contamination of benchmark datasets refers to the presence of test data in either the pre-training or fine-tuning data of an LLM. Such test data can either be present in full or as a subset, and may also include derivatives or transformed versions of the original test data. For instance, certain multilingual benchmarks like XNLI Conneau et al. (2018) and MLQA Lewis et al. (2020) are translations of a benchmark originally created in English. It is possible that the original English benchmark may be present in an LLM’s data, even if the multilingual version is not. Due to this, despite no contamination of the multilingual dataset, the LLM may still be able to recover an answer from the contaminated data through cross-lingual transfer.
Balloccu et al. (2024) introduce the concept of indirect data leakage, which happens when benchmark data is leaked when researchers evaluate models via APIs or interfaces, where the usage policy states that user interaction data can be used for improving the models. While there are access mechanisms to these models that guarantee that user data is not used for fine-tuning, such as in the case of Azure OpenAI 111https://learn.microsoft.com/en-us/azure/ai-services/openai/, it is likely that researchers performing evaluation do not use LLMs via such access mechanisms. Balloccu et al. (2024) review several papers on benchmarking and find that 42% of papers they review may have inadvertently leaked data to GPT-3.5 and GPT-4 due to their access mechanism, which does not guarantee that conversational data will not be used for fine tuning or improving models. According to the study, based on the benchmarks used in the surveyed papers, a total of 4.7M benchmark samples across 263 benchmarks have been leaked in this way.
@@
Many prompting-based techniques have been used to detect contamination by asking the model to complete model cards or data points in the dataset Ahuja et al. (2023a); Golchin and Surdeanu (2023b). Deng et al. (2023) prompt models to fill in a masked answer, which is the incorrect answer in a multiple choice setup, and also prompt the model to predict an unlikely word in a test example. The fact that open and closed LLMs are able to do this on popular benchmarks indicates contamination. Li and Flanigan (2023) use training data inspection, task example extraction, and a membership inference attack to suggest that LLM improvements can be explained by test data contamination.
Golchin and Surdeanu (2023a) propose a technique in which they generate perturbations of the benchmark datasets while preserving meaning. They provide the perturbations with the original text as options for the model to pick from, the idea being that a model contaminated with the benchmark will pick the original data point more often that the other data points even though they share the same semantics. While this approach has been somewhat effective in detecting contamination, it may not perform well in scenarios where the perturbations, generated by an LLM, are of inferior quality and do not preserve semantics, such as multilingual settings Ahuja et al. (2023b) where the LLM does not perform this task well in non-English languages.
The academic community and commercial entities such as HuggingFace and Meta are currently advocating for the use of open models to prevent many of the issues related to language models. While it is desirable for many reasons to open-source LLMs, it is likely that some models will not be made available to the public. Additionally, many open-source models do not disclose information about the training dataset, making it challenging to detect contamination. Moreover, several open-source models are constructed on top of existing open-source models which may not reveal training dataset details (e.g., fine-tuned versions of the Llama Touvron et al. (2023) models), meaning that contamination continues to be a significant challenge.
Overall, while it is an important endeavour to be able to detect contamination in models, existing techniques are not fool-proof, and hence, it is critical to prevent contamination in the first place.
@@ 3 Model Evaluation on Private Benchmarks
To prevent contamination, we propose a solution in which the benchmark dataset is kept private, which would by design, ensure that the model has never seen the benchmark dataset. However, doing so is not straightforward and faces the following two challenges: 1) model publishers of proprietary LLMs typically wish to keep their models (or at least model weights) private; how do you then evaluate models on private benchmarks while keeping the model weights private? 2) If the benchmarks are kept private, how can we audit the benchmark to ensure that it is of good quality?
We describe various scenarios in which we trust various actors in the process - the model owner, dataset owner, a trusted third party etc. We articulate different methods in which private benchmarking can be achieved in all these scenarios highlighting their pros and cons. To answer the first question (evaluating models on private benchmarks keeping model weights private), we discuss existing solutions to the problem in Section 3.1 and present recent technological advancements that can provide better solutions to this in Section 3.2. In Section 4, we answer the second question and discuss how private benchmarks can be audited.
Let us fix some notation. Let the model architecture (function description) without the model weights be denoted by . Let the weights of the model be denoted by and let the inputs in the benchmark dataset be denoted by . These inputs may also be labelled with labels . Typically the objective of an evaluation would be to evaluate every input on the model with weights and check if the output of this matches the label . That is, we want to output the bit which is when and otherwise. These results can be then collated to compute, for example, the accuracy of the model . Other functions can also be computed on the model’s outputs.
@@ 3.1.1 Model as an API (Trusted model owner)
If the model is only available as an API (e.g., in the case of proprietary models such as the OpenAI models), then one can only evaluate the benchmarks on these models through queries to the model. Since the model publisher (e.g., OpenAI) will get to see the queries being made to the model (which would be the benchmark dataset in this case), in this scenario, the model publisher would be trusted to not train/contaminate the models on the benchmarks. While in this solution, the model architecture as well as weights are kept private, the benchmark dataset is revealed to the model publisher and one must rely on the publisher ensuring that the model has not been contaminated by the benchmark. The solution architecture for this case is illustrated in Figure 1.
@@ 3.1.2 Open-source Models (Trusted dataset owner)
In this scenario, the model is open-source (e.g., Llama2 Touvron et al. (2023)) but the benchmark dataset is kept private by the benchmark owner. In this case, the model evaluation can be done on the private benchmark dataset and the results of the evaluation can be made available. The owner of the benchmark dataset performs the evaluation by hosting the model themselves and must be trusted to perform the evaluation correctly. This guarantees that the benchmark dataset does not flow back into the model for making model improvements. The solution architecture for this case is illustrated in Figure 2.
@@ 3.1.3 Trusted Evaluation of models by a Third Party (Trusted third party)
In this scenario, a trusted third party can host a computation environment that is trusted by all entities. This environment can receive both the model as well as the benchmark dataset and the evaluation of the model can be done in the clear within this environment. Hence, one relies on the trusted third party to keep both the model architecture/weights private as well as to ensure that the model has not been contaminated by the benchmark. The solution architecture for this case is illustrated in Figure 3.
@@ 3.2.1 Confidential Computation of models on the platform (Hub and Spoke)
In this scenario, the computation environment hosted will be a trusted execution environment (TEE) such as Azure Confidential Computing 222https://azure.microsoft.com/en-us/solutions/confidential-compute/, Google Confidential Computing 333https://cloud.google.com/security/products/confidential-computing or AWS Nitro System 444https://aws.amazon.com/ec2/nitro/. This TEE cannot be accessed by any entity (including the service hosting it). The TEE can be a confidential Virtual Machine (VM), a hardware enclave (such as Intel SGX) or NVIDIA H100 GPUs. The model publisher and benchmark dataset owner can encrypt the model and benchmark respectively under the public key of the confidential environment. Within the TEE, the model and the benchmark can be decrypted, evaluation can be performed, and the results of the evaluation can be made available. This solution has the advantage that both the model weights as well as the model architecture can be kept hidden from the benchmark owner. The solution architecture for this case is illustrated in Figure 4.

### v2

# : Private Benchmarking to Prevent Contamination and Improve Comparative Evaluation of LLMs

@@ Abstract
Benchmarking is the de-facto standard for evaluating LLMs, due to its speed, replicability and low cost. However, recent work has pointed out that the majority of the open source benchmarks available today have been contaminated or leaked into LLMs, meaning that LLMs have access to test data during pretraining and/or fine-tuning. This raises serious concerns about the validity of benchmarking studies conducted so far and the future of evaluation using benchmarks. To solve this problem, we propose Private Benchmarking, a solution where test datasets are kept private and models are evaluated without revealing the test data to the model. We describe various scenarios (depending on the trust placed on model owners or dataset owners), and present solutions to avoid data contamination using private benchmarking. For scenarios where the model weights need to be kept private, we describe solutions from confidential computing and cryptography that can aid in private benchmarking. We build an end-to-end system, , that enables such private benchmarking showing that the overheads introduced to protect models and benchmark are negligible (in the case of confidential computing) and tractable (when cryptographic security is required). Finally, we also discuss solutions to the problem of benchmark dataset auditing, to ensure that private benchmarks are of sufficiently high quality.
@@ 1 Introduction
Large Language Models (LLMs) have become increasingly popular due to their impressive performance in various tasks and domains, often surpassing human capabilities. However, it is crucial for LLMs to not only excel in their tasks but also be grounded in factual information and avoid generating harmful content. Therefore, evaluating and understanding the limitations and capabilities of LLMs has become an important topic of research.
@@
LLMs are commonly evaluated through automated techniques like benchmarking, or by human interaction and feedback. Holistic evaluation of LLMs [liang2022holistic] evaluates LLMs on several dimensions such as accuracy, calibration, bias and toxicity. Several popular benchmarks have been proposed such as BigBench [srivastava2023beyond] to increase the task and language coverage of benchmarking. However, benchmarking has a number of limitations. For instance, the tasks and datasets provided in standard open source benchmarks may not always be applicable in real-world scenarios. Additionally, certain benchmarks may be of low quality due to inadequate or biased instructions [parmar2023don] or limited training of annotators [geva2019we]. Moreover, some benchmarks may contain unintended artifacts introduced by annotatorsβ subconscious biases [poliak2018hypothesis]. Nevertheless, in spite of these constraints, benchmarking remains a fast, cost-effective, and replicable means of comparing multiple models or versions of a single model. As such, benchmarking has become the de-facto standard for evaluating the capabilities of LLMs, showing superiority of one LLM over another and deciding which version of a model to ship in production.
@@
However, the issue of test dataset contamination or leakage has raised significant concerns about the reliability of benchmarking and the validity of its outcomes. Benchmark datasets are commonly found on the internet, allowing large web crawlers that collect LLM pretraining data to potentially incorporate them into pretraining data. Another concerning issue, as pointed out by [balloccu2024leak], is that LLM researchers could have inadvertently leaked benchmark datasets to closed source models during LLM evaluation, and this data can be used for fine-tuning or prompt optimization. According to estimates based on their survey, over 4 million data points have been leaked to closed models such as ChatGPT. Creating benchmarks can be a costly process due to the human effort and expertise required for data creation and annotation. Therefore, it is imperative to prevent contamination of benchmarks.
@@
In this paper, we introduce the notion of Private Benchmarking - a solution that keeps the benchmark dataset private from the model, while revealing only the results of the evaluation. This ensures that the model never sees the test dataset. There are several possible scenarios - the model may be open source, in which case it can be hosted locally and a private benchmark can be run on it. In case of closed models, we may trust the model owner not to use any benchmark dataset for fine-tuning or model improvements. However, as pointed out by [balloccu2024leak], it can be challenging for model owners to identify benchmark data and filter it out to avoid training on it. In case of closed models, the model owner may want to keep models and their weights private. To account for all of the above scenarios, we build an end-to-end system, 111, stands for TRusted UnContaminated Evaluation, that implements private benchmarking in a variety of trust models (ranging from when some entity such as the dataset owner, model owner or a third party may be trusted, to scenarios where no entity may be trusted and confidential computing or cryptographic solutions must be deployed). We demonstrate the practical feasibility of our system through experimental evaluation.
Another question that becomes relevant if the benchmark datasets are kept private is: how does one guarantee that it has test data points that are of good quality? Ideally, we want solutions where an auditor is able to check a private benchmark for quality. In this work, we present unique solutions to these problems by introducing advances in confidential computing and cryptographic protocols to the NLP benchmarking community. As far as we are aware, this is the first work that has approached the problem of benchmark data contamination with this perspective.
@@ 2 Challenges with LLM benchmarking today
2.1 Lack of access to proprietary benchmarks
Major developers of LLMs continuously create benchmarks to internally evaluate and iteratively improve their models. Creating such benchmarks requires considerable expertise and time. Such internal benchmarks also hold a competitive advantage to the model owner, as their secret sauce that results in their models outperforming others. Hence model builders in the industry do not always share their benchmarks. Several commercial organizations building applications on top of LLMs also utilize their own internal benchmarks based on user data to evaluate their own models before shipping to production - however, such datasets are not typically made available for research use due to concerns about privacy and sensitivity of the data. If there are mechanisms for competing models to be evaluated on competing benchmarks with neither the model details nor the benchmarks data set being revealed to either party, it will greatly enhance credible benchmarking and faster improvements of models themselves. Access to more benchmarks created by industry will also boost LLM and evaluation research, particularly in the academic community.
@@ 2.2 Data contamination in LLMs
In the context of Large Language Models (LLMs), contamination of benchmark datasets refers to the presence of test data in either the pre-training or fine-tuning data of an LLM. Such test data can either be present in full or as a subset, and may also include derivatives or transformed versions of the original test data. For instance, certain multilingual benchmarks like XNLI [conneau2018xnli] and MLQA [lewis2020mlqa] are translations of a benchmark originally created in English. It is possible that the original English benchmark may be present in an LLMβs data, even if the multilingual version is not. Due to this, despite no contamination of the multilingual dataset, the LLM may still be able to recover an answer from the contaminated data through cross-lingual transfer. [balloccu2024leak] introduce the concept of indirect data leakage, which happens when benchmark data is leaked when researchers evaluate models via APIs or interfaces, where the usage policy states that user interaction data can be used for improving the models. [balloccu2024leak] review several papers on benchmarking and find that 42% of papers they review may have inadvertently leaked data to GPT-3.5 and GPT-4 due to their access mechanism, which does not guarantee that conversational data will not be used for fine tuning or improving models. According to the study, based on the benchmarks used in the surveyed papers, a total of 4.7M benchmark samples across 263 benchmarks have been leaked in this way.
@@
Prompting-based techniques detect contamination by asking the model to complete model cards or data points in the dataset [ahuja-etal-2023-mega, golchin2023time]. [deng2023investigating] prompt models to fill in a masked answer, which is the incorrect answer in a multiple choice setup, and also prompt the model to predict an unlikely word in a test example. [li2023task] use training data inspection, task example extraction, and a membership inference attack to suggest that LLM improvements can be explained by test data contamination. [golchin2023data] propose a technique in which they generate perturbations of the benchmark datasets while preserving meaning. They provide the perturbations with the original text as options for the model to pick from, the idea being that a model contaminated with the benchmark will pick the original data point more often that the other data points even though they share the same semantics. While this approach has been somewhat effective in detecting contamination, it may not perform well in scenarios where the perturbations, generated by an LLM, are of inferior quality and do not preserve semantics, such as multilingual settings [ahuja2023megaverse] where the LLM does not perform this task well in non-English languages.
While it is desirable for many reasons to open-source LLMs, it is likely that some models will not be made available to the public. Additionally, many open-source models do not disclose information about the training dataset, making it challenging to detect contamination. Moreover, several open-source models are constructed on top of existing open-source models which may not reveal training dataset details (e.g., fine-tuned versions of the Llama [touvron2023llama] models), meaning that contamination continues to be a significant challenge. Overall, while it is an important endeavour to be able to detect contamination in models, existing techniques are not fool-proof, and hence, it is critical to prevent contamination in the first place.
@@ 3 Model Evaluation on Private Benchmarks
To prevent contamination, we propose a solution in which the benchmark dataset is kept private, which would by design, ensure that the model has never seen the benchmark dataset. However, doing so is not straightforward and faces the following two challenges: 1) model publishers of proprietary LLMs typically wish to keep their models (or at least model weights) private; how do you then evaluate models on private benchmarks while keeping the model weights private? 2) If the benchmarks are kept private, how can we audit the benchmark to ensure that it is of sufficiently high quality?
We describe various scenarios in which we trust various actors in the process - the model owner, dataset owner, a trusted third party etc. We articulate different methods in which private benchmarking can be achieved in all these scenarios highlighting their pros and cons. To answer the first question (evaluating models on private benchmarks keeping model weights private), we discuss existing solutions to the problem in SectionΒ 3.1 and present recent technological advancements that provide better solutions to this in SectionΒ 3.2. We demonstrate the practical feasibility of our approach through an end-to-end system described and evaluated in SectionΒ 4. In SectionΒ 5, we answer the second question and discuss how private benchmarks can be audited.
Let us fix some notation. Let the model architecture (function description) without the model weights be denoted by . Let the weights of the model be denoted by and let the inputs in the benchmark dataset be denoted by . These inputs may also be labelled with labels . Typically the objective of an evaluation would be to evaluate every input on the model with weights and check if the output of this matches the label . That is, we want to output the bit which is when and otherwise. These results can be then collated to compute, for example, the accuracy of the model . Other functions can also be computed on the modelβs outputs.
@@ 3.1.1 Model as an API (Trusted model owner)
If the model is only available as an API (e.g., in the case of proprietary models such as the OpenAI models), then one can only query these models on benchmark inputs and compute the accuracy of the returned results. Since the model publisher (e.g., OpenAI) will get to see the queries being made to the model (which would be the benchmark dataset in this case), in this scenario, the model publisher would have to be trusted to not train/contaminate the models on the benchmarks. While in this solution, the model architecture as well as weights are kept private, the benchmark inputs are revealed to the model publisher and one must rely on the publisher ensuring that the model has not been contaminated by the benchmark. The solution architecture for this case is illustrated in FigureΒ 2.
(Trusted Model Owner)
(Trusted dataset owner)
@@ 3.1.2 Open-source Models (Trusted dataset owner)
In this scenario, the model is open-source (e.g., Llama2 [touvron2023llama]) but the benchmark dataset is kept private by the benchmark owner (e.g., ScaleΒ [scale]). In this case, the model evaluation can be done on the private benchmark dataset and the results of the evaluation can be made available. The owner of the benchmark dataset performs the evaluation by hosting the model themselves and must be trusted to perform the evaluation correctly. This guarantees that the benchmark dataset does not flow back into the model for making model improvements. The solution architecture for this case is illustrated in FigureΒ 2.
@@ 3.1.3 Trusted Evaluation of models by a Third Party (Trusted third party)
In this scenario, a trusted third party can host a computation environment that is trusted by all entities. This environment can receive both the model as well as the benchmark dataset and the evaluation of the model can be done in the clear within this environment. Hence, one relies on the trusted third party to keep both the model architecture/weights private as well as to ensure that the model has not been contaminated by the benchmark. The solution architecture for this case is illustrated in FigureΒ 4.
@@ 3.2 Proposed New Solutions
We propose two approaches to private benchmarking based on different technologies; seeΒ [queue] for a discussion of tradeoffs among them.
@@ 3.2.1 Confidential Computation of models on the platform (Hub and Spoke)
In this scenario, the computation environment hosted will be a trusted execution environment (TEE) such as Azure Confidential Computing 222https://azure.microsoft.com/en-us/solutions/confidential-compute/, Google Confidential Computing 333https://cloud.google.com/security/products/confidential-computing or AWS Nitro System 444https://aws.amazon.com/ec2/nitro/. This TEE cannot be accessed by any entity (including the service hosting it). The TEE can be a confidential Virtual Machine (VM), a hardware enclave (such as Intel SGX) or Hopper and Blackwell GPUs by NVIDIA in confidential mode. The model publisher and benchmark dataset owner can encrypt the model and benchmark respectively under the public key of the confidential environment. Within the TEE, the model and the benchmark can be decrypted, evaluation can be performed, and the results of the evaluation can be made available. This solution has the advantage that both the model weights as well as the model architecture can be kept hidden from the benchmark owner. The solution architecture for this case is illustrated in FigureΒ 4. While confidential computing had scalability issues in the past because of the small size of the memory in the TEE, this is no longer an issue in the latest systems based on confidential GPUsΒ [hoppercc].
(Trusted Third Party)
(Hub and Spoke)